# FIZZ: Factual Inconsistency Detection by Zoom-in Summary and Zoom-out Document

### v2

@@ 2 Related Work
SummaC Laban et al. (2022) re-visit NLI-based models and granularity choice for inconsistency detection in summarization.  AlignScore Zha et al. (2023) develops an alignment system, incorporating a summarization consistency checking metric and an NLI model, which has been trained across a diverse array of tasks that can be aligned with NLI. Our proposed FIZZ is also based on NLI. However, unlike the aforementioned systems, which compare the summary at the sentence level, FIZZ conducts comparisons at a more fine-grained atomic fact level with high interpretability.
@@ 3 FIZZ
The overall flow of our proposed system FIZZ is presented in Figure 2. Our method first begins with the application of a coreference resolution model to a given $(document,summary)$ pair, resulting in a new pair of texts $(document,summary)$ where coreferences have been resolved (Section 3.1). Following this, we proceed to generate atomic facts from the coreference-resolved summary leveraging LLMs as a zooming-in approach for the summary. (Section 3.2). Using the generated atomic facts, we compute the score of each atomic fact with the NLI system. (Section 3.3). Finally, we propose a granularity expansion method, which is a way of zooming out the documents, to compute the score for the summaries that contain high abstractiveness more accurately.
@@ 4.1 Experimental Setups
In our experiments, we leverage MT5 Bohnet et al. (2023) for coreference resolution, which returns with the identification of clusters referring to the same entities. With these clusters, we further implement rule-based pronoun substitution strategies to generate coreference resolved texts. For atomic fact decomposition, the Orca-2 model Mitra et al. (2023) is utilized. Additionally, this work adopts the same off-the-shelf NLI model as implemented in SummaC (See Appendix B for more details).
@@ 4.3 Baselines
AggreFact-Cnn AggreFact-XSum
FtSota EXf Old FtSota EXf Old Avg
Baseline 50.0 50.0 50.0 50.0 50.0 50.0 50.0
DAE* 59.4 67.9 69.7 73.1 - - 67.5
QuestEval 63.7 64.3 65.2 61.6 60.1 59.7 62.4
SummaC-ZS 63.3 76.5 76.3 56.1 51.4 53.3 62.8
SummaC-Cv 70.3 69.8 78.9 67.0 64.6 67.5 69.7
QAFactEval 61.6 69.1 80.3 65.9 59.6 60.5 66.2
AlignScore 53.4 73.1 80.2 70.2 80.1 63.7 70.1
ChatGPT-ZS 66.2 64.5 74.3 62.6 69.2 60.1 66.2
ChatGPT-CoT 49.7 60.4 66.7 56.0 60.9 50.1 57.3
ChatGPT-DA 48.0 63.6 71.0 53.6 65.6 61.5 60.6
ChatGPT-Star 55.8 65.8 71.2 57.7 70.6 53.8 62.5
FactScore 69.9 71.6 73.9 68.0 63.5 66.8 69.0
FacTool 72.7 66.1 60.8 68.0 64.0 62.2 65.6
FIZZ 73.2 67.3 76.0 69.7 72.4 68.5 71.2
${w/o\,\,GE}$ 68.0 67.1 77.6 69.4 71.5 69.8 70.6
${w/o\,\,AF}$ 72.4 67.0 80.4 64.9 67.1 66.1 69.7
Table 1: Balanced accuracy results on the AggreFact dataset. Following  Tang et al. (2023), we omitted the results from DAE, as it was trained on the XSumFaith Goyal and Durrett (2021) dataset, which includes human-annotated summaries from EXformer and Old. Highest performance is highlited in bold, and the second highest is underlined.
@@
Table 2: Balanced accuracy results using a single threshold with 95% confidence intervals on the AggreFact-FtSota split dataset.
Additionally, we incorporate FActScore Min et al. (2023) and FacTool Chern et al. (2023) in our baselines. These methods decompose generated texts into atomic facts and then retrieve corresponding entries from a given knowledge base, such as Wikipedia, to evaluate the factuality of the generated context. For the purpose of verification, we assume the availability of this knowledge base, which we use as the source document to assess summary factual consistency. In FActScore, we employ a No-context LM for factual verification. This approach operates on a QA basis, assessing whether atomic facts are true or false with respect to the source document. In FacTool, we utilize a Knowledge-based QA approach. This also follows a QA format but incorporates the CoT method, where the LLM evaluates if claims are true or false relative to the source document [^2] Details of the experiments are provided in Appendix B.
@@ 4.5 Analysis
To investigate the most suitable LLMs for generating atomic facts, we evaluate the generation of atomic facts using various LLMs, including gpt-3.5-turbo, gpt-3.5-turbo-instruct, and other 7B models such as Zephyr Tunstall et al. (2023) and Mistral Jiang et al. (2023) [^3] Details of the model are provided in Appendix B. The results, documented in Table 3, demonstrate that while the atomic facts generated by gpt-3.5-turbo and gpt-3.5-turbo-instruct generally perform better compared to other metrics, they are still inferior to those produced by Orca-2. The performance drop associated with the gpt series suggests a noteworthy observation. We explain that this discrepancy is due to the length of the atomic facts. As shown in Table 3, which includes the average token length of atomic facts after the filtering process per summary, there is a clear inverse relationship between the number of tokens in an atomic fact and its average performance. Longer atomic facts tend to contain more entities and are less concise. Such sentences are less suitable as hypotheses when compared sentence-wise using NLI models. Furthermore, the sensitivity of using the minimum atomic fact scores as the final score exacerbates the challenge, making it difficult to achieve desired outcomes with lengthy sentences. In contrast, other 7B models such as LLaMa Touvron et al. (2023) show limitations in adhering to instructions for atomic fact decomposition.
Sizes of Granularity Expansion
@@
Despite the increase in average score for the maximum of four sentences, the scores for CNN summaries actually declined. Additionally, we observe that computational costs rose with increasing granularity. Hence, we determined that the maximum of three sentences represents the best trade-off between computational cost and performance.
@@
Table 4: Size of granularity choice in granularity expansion on AggreFact-FtSota split. We tested four granularity sizes on the document side for the additional lookup of an atomic fact which scored too low entailment score. s/it indicates seconds per iteration during the inference of an NLI model.
Effectiveness of Coreference Resolution
Moreover, Table 5 presents more granular experimental results on the impact of coreference resolution. We implemented experiments to evaluate the impact of coreference resolution on both documents and atomic facts. Our investigation included scenarios where coreference resolution was applied and cases where it was not. We show that texts with resolved coreferences, whether they be atomic facts or documents, consistently outperform those without resolution. Notably, there is a marked improvement in performance on datasets based on CNN Hermann et al. (2015) summaries compared to those based on XSum summaries. This is likely due to the extractive nature of CNN-based summaries, as opposed to the more abstractive summaries derived from XSum.
@@
Table 5: Effect of coreference resolution of document and atomic facts on AggreFact-FtSota splits before the process of granularity expansion. Coreference resolved either atomic fact or document both show an increase in the average.
@@ 5 Conclusion
This work uses English document summarization dataset, AggreFact. This dataset is publicly available online. We also provided adequate citations for the papers and sources we consulted in writing our paper. As our work addresses consistency checking, Our work may have implications for society in terms of preventing the spread of inaccurate information, as it deals with factual consistency checking.
@@ Appendix A Prompt for Atomic Facts Decomposition
Appendix B Experimental Details
Baseline Details
@@ Appendix B Experimental Details
Language Model Details
@@
NLI Model Details
We leveraged NLI model from SummaC [^9] https://huggingface.co/tals/albert-xlarge-vitaminc-mnli, which has been trained on the conventional NLI datasets SNLI Bowman et al. (2015), MNLI Williams et al. (2018), ANLI Nie et al. (2020), and also on VitaminC Schuster et al. (2021).
Coreference Resolution Details
@@
Other Details

### v3

@@ 2 Related Work
SummaC Laban et al. (2022) re-visit NLI-based models and granularity choice for inconsistency detection in summarization.  AlignScore Zha et al. (2023) develops an alignment system, incorporating a summarization consistency checking metric and an NLI model, which has been trained across a diverse array of tasks that can be aligned with NLI. The recently proposed method, FENICE Scirè et al. (2024), also aligns decomposed atomic facts with several document sentences, but it lacks interpretability on summary side. Our proposed system, FIZZ, is also based on NLI. However, unlike the aforementioned systems, which mostly compare the summary at the sentence level, FIZZ conducts comparisons at a more fine-grained atomic fact level with high interpretability.
@@ 3 FIZZ
The overall flow of our proposed system FIZZ is presented in Figure 2. Our method first begins with the application of a coreference resolution model to a given $(document,summary)$ pair, resulting in a new pair of texts $(document,summary)$ where coreferences have been resolved (Section 3.1). Following this, we proceed to generate atomic facts from the coreference-resolved summary leveraging LLMs as a zooming-in approach for the summary (Section 3.2). Using the generated atomic facts, we compute the score of each atomic fact with the NLI system (Section 3.3). Finally, we propose a granularity expansion method, which is a way of zooming out the documents, to compute the score for the summaries that contain high abstractiveness more accurately.
@@ 4.1 Experimental Setups
In our experiments, we leverage MT5 Bohnet et al. (2023) for coreference resolution, which returns with the identification of clusters referring to the same entities. With these clusters, we further implement rule-based pronoun substitution strategies to generate coreference resolved texts. For atomic fact decomposition, the Orca-2 model Mitra et al. (2023) is utilized. Additionally, this work adopts the same off-the-shelf NLI model as implemented in SummaC (See Appendix D for more details).
@@ 4.3 Baselines
Table 1: Balanced accuracy using a single threshold with 95% confidence intervals on the AggreFact-FtSota split dataset. Highest performance is highlited in bold, and the second highest is underlined.
Additionally, we incorporate FActScore Min et al. (2023) and FacTool Chern et al. (2023) in our baselines. These methods decompose generated texts into atomic facts and then retrieve corresponding entries from a given knowledge base, such as Wikipedia, to evaluate the factuality of the generated context. For the purpose of verification, we assume the availability of this knowledge base, which we use as the source document to assess summary factual consistency. In FActScore, we employ a No-context LM for factual verification. This approach operates on a QA basis, assessing whether atomic facts are true or false with respect to the source document. In FacTool, we utilize a Knowledge-based QA approach. This also follows a QA format but incorporates the CoT method, where the LLM evaluates if claims are true or false relative to the source document. Details of the experiments are provided in Appendix B.
AggreFact-Cnn AggreFact-XSum
FtSota EXf Old FtSota EXf Old Avg
Baseline 50.0 50.0 50.0 50.0 50.0 50.0 50.0
DAE* 59.4 67.9 69.7 73.1 - - 67.5
QuestEval 63.7 64.3 65.2 61.6 60.1 59.7 62.4
SummaC-ZS 63.3 76.5 76.3 56.1 51.4 53.3 62.8
SummaC-Cv 70.3 69.8 78.9 67.0 64.6 67.5 69.7
QAFactEval 61.6 69.1 80.3 65.9 59.6 60.5 66.2
AlignScore 53.4 73.1 80.2 70.2 80.1 63.7 70.1
ChatGPT-ZS 66.2 64.5 74.3 62.6 69.2 60.1 66.2
ChatGPT-CoT 49.7 60.4 66.7 56.0 60.9 50.1 57.3
ChatGPT-DA 48.0 63.6 71.0 53.6 65.6 61.5 60.6
ChatGPT-Star 55.8 65.8 71.2 57.7 70.6 53.8 62.5
FactScore 69.9 71.6 73.9 68.0 63.5 66.8 69.0
FacTool 72.7 66.1 60.8 68.0 64.0 62.2 65.6
FIZZ (Ours) 73.2 67.3 76.0 69.7 72.4 68.5 71.2
Table 2: Balanced accuracy on the AggreFact dataset. As in Tang et al. (2023), we omitted the results from DAE, as it was trained on the XSumFaith Goyal and Durrett (2021) dataset, which includes human-annotated summaries from EXformer and Old.
@@ 4.4 Results
ROUGE-1 Avg. Number of Avg. Token
P R F1 Atomic Facts Length
Human 1.00 1.00 1.00 8.7 98.4
Orca-2 0.70 0.69 0.68 8.7 96.3
gpt-3.5-turbo 0.78 0.84 0.79 7.8 105.0
gpt-3.5-turbo-instruct 0.73 0.72 0.70 13.0 149.6
Mistral 0.63 0.62 0.61 9.6 104.1
Zephyr 0.51 0.60 0.52 10.1 122.0
Table 4: Experimental results of generated atomic facts on RoSE dataset. The results with the highest human correlation are highlighted in bold.
@@ 4.5 Analysis
To investigate the most suitable LLMs for generating atomic facts, we evaluate the generation of atomic facts using various LLMs, including gpt-3.5-turbo, gpt-3.5-turbo-instruct, and other 7B models such as Zephyr Tunstall et al. (2023) and Mistral Jiang et al. (2023). The results, documented in Table 3, demonstrate that while the atomic facts generated by gpt-3.5-turbo and gpt-3.5-turbo-instruct generally perform better compared to other metrics, they are still inferior to those produced by Orca-2. The performance drop associated with the gpt series suggests a noteworthy observation. We explain that this discrepancy is due to the length of the atomic facts. As shown in Table 3, which includes the average token length of atomic facts after the filtering process per summary, there is a clear inverse relationship between the number of tokens in an atomic fact and its average performance. Longer atomic facts tend to contain more entities and are less concise. Such sentences are less suitable as hypotheses when compared sentence-wise using NLI models. Furthermore, the sensitivity of using the minimum atomic fact scores as the final score exacerbates the challenge, making it difficult to achieve desired outcomes with lengthy sentences. In contrast, other 7B models such as LLaMa Touvron et al. (2023) show limitations in adhering to instructions for atomic fact decomposition. Details of the model usage are provided in Appendix C.
@@
In previous studies Zhang and Bansal (2021); Chern et al. (2023); Scirè et al. (2024), the evaluation of the quality and the completeness of the LLM generated atomic facts focuses solely on content similarity (i.e., ROUGE-1) with human-written atomic facts. However, we consider content similarity evaluation to be insufficient and added two additional factors: 1) Average token length in atomic facts and 2) Average number of atomic facts. In Table 3, we demonstrate the correlation between the average token length of atomic facts and overall performance. Building on this, we now analyze the token length of both human-written and generated atomic facts. Additionally, since the content similarity metric does not take into account the number of atomic facts, we also include the average number of atomic facts in our results. We report the comparative analysis of the LLM generated atomic facts against human-written atomic facts in Table 4. The experiments were implemented using the RoSE Liu et al. (2023) dataset, which includes 2,500 summaries and their corresponding human-written atomic facts. As shown in the experimental results, gpt-3.5-turbo demonstrates the highest capability by achieving the top score in content similarity. However, it shows a significant difference in the number of atomic facts and the number of tokens in atomic facts. In contrast, Mistral scores lower in content similarity but exhibits higher human correlation in the number of atomic facts and token lengths. The model that achieves the highest human correlation in both the number of atomic facts and token lengths is Orca-2, which shows the best performance among LLMs as in Table 3. These findings suggest that while content similarity is important, the number of atomic facts and token lengths are equally critical factors to consider. Details on computing content similarity are provided in Appendix G.
@@
Table 5: Size of granularity choice in granularity expansion on AggreFact-FtSota split. s/it indicates seconds per iteration for the inference of an NLI model.
Sizes of Granularity Expansion
Despite the increase in average score for the maximum of four sentences, the scores for CNN summaries actually declined. Additionally, we observe that computational costs rose with increasing granularity. Hence, we determined that the maximum of three sentences represents the best trade-off between computational cost and performance. Details on granularity expansion condition choice are provided in Appendix F.
@@
Table 6: Effect of coreference resolution of document and atomic facts on AggreFact-FtSota splits before the process of granularity expansion.
Effectiveness of Coreference Resolution
Moreover, Table 6 presents more granular experimental results on the impact of coreference resolution. We implemented experiments to evaluate the impact of coreference resolution on both documents and atomic facts. Our investigation included scenarios where coreference resolution was applied and cases where it was not. We show that texts with resolved coreferences, whether they be atomic facts or documents, consistently outperform those without resolution. Notably, there is a marked improvement in performance on datasets based on CNN Hermann et al. (2015) summaries compared to those based on XSum summaries. This is likely due to the extractive nature of CNN-based summaries, as opposed to the more abstractive summaries derived from XSum. Details on coreference resolution usage are provided in Appendix E.
@@ 5 Conclusion
This work uses English document summarization dataset, AggreFact. This dataset is publicly available online. We also provided adequate citations for the papers and sources we consulted in writing our paper. Our work may have implications for society in terms of preventing the spread of inaccurate information, as it deals with factual consistency checking.
@@ Appendix A Prompt for Atomic Facts Decomposition
Appendix B Details on Baselines
@@
Appendix C Details on the Usage of Large Language Models
@@
Appendix D Details on the Usage of NLI Model
In this study, we tried to analyze the effect of our proposed atomic fact level decomposition instead of using entire sentences. To ensure a fair comparison of our approach with SummaC, which demonstrated the best performance using whole sentences, we employed the same NLI model that was utilized in SummaC [^7] https://huggingface.co/tals/albert-xlarge-vitaminc-mnli. The model has been trained on the conventional NLI datasets SNLI Bowman et al. (2015), MNLI Williams et al. (2018), ANLI Nie et al. (2020), and also on VitaminC Schuster et al. (2021).
In Table 7, we present the performance results of various NLI models. Specifically, we have included the results for DeBERTa-large-mnli [^8] https://huggingface.co/MoritzLaurer/DeBERTa-v3-large-mnli-fever-anli-ling-wanli and RoBERTa-large-pyrxsum [^9] https://huggingface.co/shiyue/roberta-large-pyrxsum. The average performance scores for DeBERTa and RoBERTa are 68.7 and 68.5, respectively. Although these scores are lower than that of ALBERT, they surpass the previous best score of 67.8 achieved by DAE on the FtSota split.
NLI Model
AggreFact-
Cnn-FtSota
AggreFact-
XSum-FtSota
Avg
ALBERT 72.6$\pm$3.0 69.3$\pm$1.9 71.0
DeBERTa 67.3$\pm$3.0 70.1$\pm$1.9 68.7
RoBERTa 70.5$\pm$3.0 66.5$\pm$1.9 68.5
Table 7: Performance of different NLI models on AggreFact-FtSota split.
Appendix E Details on the Usage of Coreference Resolution
@@
Appendix F Details on Granularity Expansion
[…]
Condition
AggreFact-
Cnn-FtSota
AggreFact-
XSum-FtSota
Avg
!(e>c & e>n) 72.6$\pm$3.0 69.3$\pm$1.9 71.0
!(e>c || e>n) 71.1$\pm$2.9 68.7$\pm$1.9 69.9
Table 8: Granularity Expansion condition choice on AggreFact-FtSota split.
Appendix G Details on Computing Content Similarity
The content similarity (ROUGE-1) in Table 4 was conducted using the following equation:
$\frac{1}{N_{data}}\sum_{N_{data}}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\max_{j=1}^{N_{g}}\left(\text{ROUGE}(c_{i},g_{j})\right)$ (5)
where $c$ denotes LLM generated atomic facts and $g$ denotes human-written atomic facts.
Appendix H Other Details